# Machine Learning to Predict Spectral Anisotropy in Valence-to-Core X-ray Emission Spectroscopy

### v1

@@
Charles A. Cardot, John Tichenor, Seth M. Shjandemaar,
Josh J. Kas, Fernando D. Vila, Gerald T. Seidler, John J. Rehr
@@
Polarization analysis in x-ray spectroscopy provides an orientation dependent sensitivity to local bonding environments. For a cluster of atoms, polarization sensitivity is most often discussed through the lens of point group symmetries. However, this is a discrete, qualitative method of classifying clusters, and it does little to indicate the degree of spectral anisotropy. Here we adopt a random forest model for quantitative prediction of spectral anisotropy. Its input relies on simplified local geometric and chemical information that can be obtained from any crystal structure file. The model is trained on over 10,000 experimentally realized transition metal structures from the Materials Project, with the target being VtC-XES calculated using the real space Green’s function code FEFF. We find that the model can strongly predict the degree of spectral anisotropy, with the primary factors being derived from the spatial moments of ligands in a cluster.
@@ 1 Introduction
Anisotropic materials have a multitude of applications in a wide variety of fields, including aerospace engineering, high speed electronics, thermo-, photo-, and piezo-electronics, and more [chen_2022, gong_2017, li_2021]. These materials exhibit directional dependence in mechanical, electronic and magnetic properties, making them highly desirable for use in a wide variety of devices, including optical components and polarization detectors, high-speed transistors, lightweight load-bearing devices, and energy storage solutions to name a few. There are many techniques for measuring the anisotropy of a given property [loudon-1964, cahill-2003, ziman-1972]. However, x-ray spectroscopy and related techniques are unique in their ability to probe the anisotropy of the local atomic, magnetic, and electronic structure of the material, allowing analysis of the fundamental structure-function relationships [tortora-2024].
Polarization is prevalent in x-ray spectroscopy as a method of analyzing spin and orbital magnetic moments [van-der-laan-2014], magnetic axes or interfaces [jo-2004, van-der-laan-2011], crystal field effects [bergmann-2002], and coordination environments [jansing-2016]. With the advent and improvement of high brilliance synchrotron radiation over the last 40 years, a diverse suite of experimental capabilities have emerged that can use polarization as an additional probe within x-ray spectroscopy to extract electronic and chemical information [glatzel-2004, pollock-2015, ketenoglu-2022, trippe-2014, stohr-1992]. In x-ray absorption spectroscopy (XAS) it can be used as a spin-dependent probe of the unoccupied density of states via x-ray magnetic circular dichroism (XMCD) [stohr-1999], or as a probe of bond directions and magnetic axes in transition metal oxides through x-ray linear dichroism (XLD) [bianconi-1988].
@@
Local anisotropy is regularly treated in terms of discrete classifications, namely point group symmetries [watanabe-1966, brener-2024, nielsen-2024]. These are used in conjunction with oxidation state to interpret the ground state electron configuration for transition metal oxides in a given crystal field (ex: high-spin and low-spin) [lafuerza-2020]. However, it is difficult to predict a priori how large of an effect the point group derived crystal field will have on a general x-ray spectrum. This is illustrated in Figure 1, which shows how the polarization dependent VtC-XES changes when going from a 5% stretch of the apical O bond-length (a, b) to a 50% stretch (c, d). While the extent of broken octahedral (Oh) symmetry is vastly different, these clusters have the same point group (D4h).
@@
Thus, there is a need for a continuous, quantitative metric that encodes local anisotropic environment in a way that can be used for predicting polarization dependence. Such a metric would enable rapid interpretation of experimental data and support the development of artificial intelligence (AI) driven workflows at synchrotron facilities. AI-based tools have already been successfully applied to phase identification, as well as parameter extraction in x-ray diffraction (XRD) and x-ray reflectometry (XRR), enabling continuous and rapid data-to-decision pipelines for beamline science [pithan-2023]. This metric can also be used for bulk, quantitative screening of large materials databases to identify candidates with strong local anisotropy effects. This kind of targeted exploration speeds up the search for functional materials in areas like spintronics [zutic-2004, wang-2023, tombros-2008], quantum information [shen-2022, hirohata-2025], and energy conversion and storage [sendek-2018].
@@
Recent advances in ab initio theoretical methods have made it possible to reliably calculate a wide range of x-ray spectroscopies [vinson-2011, apra-2020]. Despite these successes, such calculations remain computationally demanding and are often limited by system size and the need to treat electron correlation [pople-1999]. By contrast, a trained machine learning model can make predictions orders of magnitude faster, enabling rapid exploration of material spaces.
@@
The relationship between a material’s local chemical and geometric environment, and its polarization-dependent spectral response, is inherently nonlinear and highly complex. Spectral anisotropy arises from directional variations in the electronic states, which are influenced by factors such as local bonding asymmetry, ligand field distortions, coordination number, and electronic hybridization [Drager.Brummer1984]. Traditional linear or descriptor-based approaches struggle to capture these coupled, high-dimensional effects. Machine learning models, by contrast, are well-suited to learn nonlinear mappings from structural and chemical features to spectral response.
@@
Machine learning has successfully been applied to several modalities of spectroscopy including x-ray absorption near-edge structure (XANESNet) [rankine-2020], extended x-ray absorption fine structure [timoshenko-2018], and vibrational spectroscopy [gastegger-2017]. The broader field of materials sciences has also demonstrated that machine learning models can successfully predict bulk material properties, such as utilizing a message-passing graph neural network (MPGNN) to capture anisotropic properties such as elastic [wen-2024] and dielectric tensors [lou-2024]. However, these GNN approaches can be limited by the scarcity of high-quality training data and, like most neural network-based methods, suffer from a black-box nature that limits interpretability.
@@
In contrast, random forest models have demonstrated strong performance in materials spectroscopy; for example, Zheng et al. successfully applied them to identify coordination environments in XANES spectra [Zheng.etal2020]. Random forests can capture nonlinear and interacting effects among structural and chemical features, making them well-suited for predicting complex spectral responses while maintaining interpretability. To the authors’ knowledge, anisotropy in the VtC-XES regime remains largely unexplored and has not yet been the subject of any machine learning applications. In this paper, we develop a random forest regression (RFR) model trained on structural and chemical descriptors of 3dd transition metal complexes to predict the spectral anisotropy in VtC-XES. To quantify the degree of anisotropy, we introduce a new metric termed the spectral anisotropy matrix sum (SAMS).
@@
The rest of this article is organized as follows. In section 2 we described the procedure for quantifying the spectral anisotropy and how we extract input features as our predictors in the model. We also address the calculation pipeline used for developing the training data set, and the procedure used to develop the random forest model. In section 3 we present the core results of the model, namely the $R^{2}$ and mean absolute error (MAE). We also assess the relative importance of the features and compare predictive capabilities across the 10 3dd transition metals. Finally in section 4 we discuss constrains of the model and propose future directions for how it could be improved and extended to new applications.
@@ 2.1 Spectral Anisotropy Matrix
To quantify the degree of anisotropy in a spectrum we introduce a new metric based on the $L^{2}$-norm or Euclidean norm of the difference between individual polarizations [bishop-2007]. We start by breaking a spectrum into the emission for three orthogonal polarization axes ($\sigma_{i},\sigma_{j},\sigma_{k}$). Then by calculating the square root of the integral over the difference squared we arrive at a single number to characterize the ”difference” between two curves. This value is normalized by the integral of the average (isotropic) spectra, $\bar{\sigma}$. We calculate this parameter for each possible pair of axes to produce the spectral anisotropy matrix (SAM) as defined in Eq. 1.
@@
The SAM is a symmetric matrix with zeroes along the diagonal. A visualization of the calculation process is shown in Figure 2, where the Cr VtC-XES is calculated along the $x$, $y$, and $z$ crystal axes of an example transition metal crystal system, LiCrCO3F2. The crystal structure is shown in (a), where the local structure around the Cr atoms is distorted such that there are no degenerate axes. The polarized spectra along the three crystal axes is shown in (b) along with the difference curves between each pair of spectra. These are then used to calculate the Euclidean norm and build the SAM shown in (c). Taking the sum of the upper off diagonal components as shown in Eq. 2, we arrive a quantitative measure, the spectral anisotropy matrix sum (SAMS). This metric captures the total amount of anisotropy in the spectrum, and is the numerical value we aim to predict with our model in this work.
A SAM is highly dependent on the basis used to define the cluster around the transition metal ion. Our goal is to create a dataset of the maximum possible SAMS signal that can be achieved for a given structure, which often requires rotating to the principal axes. To address this, we choose to operate in the basis that diagonalized the quadrupole matrix (both for calculating the SAMS and all other features used in the model), which provides a more natural basis for maximizing the spectral anisotropy. The relationship between the spectra and the quadrupole moment will be discussed in more detail in section 2.3.
@@ 2.2 Feature Anisotropy Matrices
To parameterize the anisotropy of the local atomic environment around an absorber atom, we construct 3 related but distinct metrics derived from the quadrupole moment (Eq. 3a), dipole moment (Eq. 3b), and inertia tensor (Eq. 3c) respectively.
@@
The quadrupole and inertia matrices are identical with the exception of the effective charge $q_{n}$, which serves to encode chemical information. We will see that the anisotropy metrics of both are strongly correlated with the spectral anisotropy, but we choose to use the eigenvectors of the weighted quadrupole moment to define a rotation-invariant basis. This is done both to maximize the observed spectral anisotropy in Eq. 2 and provide a set coordinate system to work in. The main reason for including both the quadrupole Eq. 3a and the inertia Eq. 3c comes from the reliance on estimating the ligand charges $q_{n}$. We rely on simple oxidation state estimations and avoid doing any direct electronic structure calculations for generating input features, and uncertainties in the charge estimates can degrade the reliability of the quadrupole metric. Incorporating a purely geometric metric thus provides a complementary, charge-independent measure of local anisotropy that can help disentangle geometric effects from uncertainties in ligand charge assignment.
The quadrupole, dipole, and inertia anisotropy matrices (denotes QAM, DAM, and IAM respectively) are shown in Eq. 4. They are all expressed in the eigen-basis of the quadrupole matrix (eigenvalues $\lambda$), where $\tilde{D}=V^{-1}DV$ and $\tilde{I}=V^{-1}IV$ are the dipole and inertia matrices transformed into this basis and $V$ is the matrix whose columns are the orthonormal eigenvectors of the weighted quadrupole moment.
@@
Consistent with the single anisotropy measure used for summarizing the spectral anisotropy metric into the SAMS in Eq. 2, a similar sum is applied to the off diagonal elements of Eq. 4 to arrive at the quadrupole, dipole, and inertia anisotropy matrix sums, denoted QAMS, DAMS, and IAMS respectively (Eq. 5). If a unit cell contains multiple unique sites for a 3dd transition metal, we average the anisotropy matrix sum over all sites to obtain a single representative number for the entire crystal.
@@ 2.3 Connection between Quadrupole and Spectra
The continuous form of the quadrupole moment tensor is defined in Eq. 7 (without additional distance weighting). This gives the quadrupole moment of a charge distribution with electron density $\rho_{e}(\vec{r})$ (Eq. 8), where $e$ is the electron charge, $N$ is the number of electrons, and $\Psi$ is the many-body wavefunction.
@@
In Eq. 6, the initial and final state wavefunctions are shaped by the local chemical environment. This manifests as polarization-dependent effects in the spectrum. Since the electron density in 8 is derived from the modulus squared of the many-body wavefunction, it retains similar spatial anisotropy, albeit without phase information. The quadrupole moment is therefore a natural option for encoding this anisotropy, capturing both geometric and electronic symmetry-breaking features of the local environment.
To further assess the relationship between the QAMS and the SAMS, we now examine how they respond to systematic geometric distortions in representative materials, specifically NiO and Cr2O3, as shown in Figure 3. Subplots (b) and (d) show the crystal structure for NiO and Cr2O3 respectively, while subplots (a) and (c) show the QAMS vs. SAMS for different degrees of stretch or compression from the experimental structure. In (a) the NiO structure is stretched along the $c$-axis in increments of 0.5%. Starting from the lower left-hand corner of the plot—where all bond lengths are equal—the structure is isotropic, and both QAMS and SAMS are exactly zero, which is in agreement with the perfect cubic NiO structure. As the structure becomes increasingly stretched, both QAMS and SAMS vary linearly. This suggests that for simple distortions such as bond elongation without significant angular rearrangement, QAMS tracks spectral anisotropy with high fidelity.
@@
However, in the case of Cr2O3 (c) we observe a distinctly nonlinear relationship between QAMS and SAMS. Here the local environment around the Cr ion is a distorted Oh, and has non-zero QAMS and SAMS for even the starting crystal structure. In the extremes of both compression and stretching, the relationship between the two parameters remains linear. In the region around the unchanged structure, we note that the relationship deviates from linearity, and that the minimum of the SAMS does not line up with the minimum of the QAMS.
This behavior leads to two key insights. First, the nonlinear behavior arises from changes not only in bond lengths but also in bond angles, which substantially affect the spatial charge distribution and, consequently, the spectral anisotropy. For NiO, the stretching occurs uniformly without altering angular geometry, but Cr2O3 has a more complex structural responses to distortion. While the two parameters are clearly correlated, their relationship is not strictly proportional. Second, it’s possible to have a local cluster that produces a zero QAMS and a non-zero SAMS (for example, a perfectly linear molecule where the ligands have the same charge q). To address this issue, we require additional features that can more effectively capture and discriminate between the sources of anisotropy. The first of these is the DAMS from Eq. 5b, which provides orthogonal information compared to the QAMS due one being derived from a metric that is odd under inversion while the other is even. A demonstration of how DAMS can help address edge cases is provided in Supplemental Information section 6.1.1

### v2

@@
Charles A. Cardot, John R. Tichenor, Seth M. Shjandemaar,
Josh J. Kas, Gerald T. Seidler, John J. Rehr
@@
Polarization analysis in x-ray spectroscopy provides an orientation dependent sensitivity to local bonding environments. For a cluster of atoms, polarization sensitivity is most often discussed through the lens of point group symmetries. However, this is a discrete, qualitative method of classifying clusters, and it does little to indicate the degree of spectral anisotropy. Here we adopt a random forest model for quantitative prediction of spectral anisotropy. Its input relies on simplified local geometric and chemical information that can be obtained from any crystal structure file. The model is trained on approximately 11,500 experimentally realized transition metal structures from the Materials Project, with the target being VtC-XES calculated using the real space Green’s function code FEFF. We find that the model can strongly predict the degree of spectral anisotropy, with the primary factors being derived from the spatial moments of ligands in a cluster.
@@ 1 Introduction
Anisotropic materials have a multitude of applications in a wide variety of fields, including aerospace engineering, high speed electronics, thermo-, photo-, and piezo-electronics, and more chen_2022, gong_2017, li_2021. These materials exhibit directional dependence in mechanical, electronic and magnetic properties, making them highly desirable for use in a wide variety of devices, including optical components in polarization detectors, high-speed transistors, lightweight load-bearing devices, and energy storage solutions. There are many techniques for measuring the anisotropy of material properties loudon-1964, cahill-2003, ziman-1972. However, x-ray spectroscopy and related techniques are unique in their ability to probe the anisotropy of the local atomic, magnetic, and electronic structure of the material, allowing analysis of the fundamental structure-function relationships tortora-2024. Polarization is prevalent in x-ray spectroscopy as a method of analyzing spin and orbital magnetic moments van-der-laan-2014, magnetic axes or interfaces jo-2004, van-der-laan-2011, crystal field effects bergmann-2002, and coordination environments jansing-2016. With the advent and improvement of high brilliance synchrotron radiation over the last 40 years, a diverse suite of experimental capabilities have emerged that can use polarization as an additional probe within x-ray spectroscopy to extract electronic and chemical information glatzel-2004, pollock-2015, ketenoglu-2022, trippe-2014, stohr-1992. In x-ray absorption spectroscopy (XAS) it can be used as a spin-dependent probe of the unoccupied density of states via x-ray magnetic circular dichroism (XMCD) stohr-1999, or as a probe of bond directions and magnetic axes in transition metal oxides through x-ray linear dichroism (XLD) bianconi-1988.
@@
Local anisotropy is regularly treated in terms of discrete classifications, namely point group symmetries watanabe-1966, brener-2024, nielsen-2024. These are used in conjunction with oxidation state to interpret the ground state electron configuration for transition metal oxides in a given crystal field (ex: high-spin and low-spin) lafuerza-2020. However, it is difficult to predict a priori how large of an effect the point group derived crystal field will have on a general x-ray spectrum. This is illustrated in Figure 1, which shows how the polarization dependent VtC-XES changes when going from a 5% stretch of the apical O bond-length Figure 1a,b to a 50% stretch Figure 1c,d. While the extent of broken octahedral (Oh) symmetry is vastly different, these clusters have the same point group (D4h).
@@
Thus, there is a need for a continuous, quantitative metric that encodes the local anisotropic environment in a way that can be used for predicting polarization dependence. Such a metric would enable rapid interpretation of experimental data and support the development of artificial intelligence (AI) driven workflows at synchrotron facilities. AI-based tools have already been successfully applied to phase identification, as well as parameter extraction in x-ray diffraction (XRD) and x-ray reflectometry (XRR), enabling continuous and rapid data-to-decision pipelines for beamline science pithan-2023. This metric can also be used for bulk, quantitative screening of large materials databases to identify candidates with strong local anisotropy effects. This kind of targeted exploration speeds up the search for functional materials in areas like spintronics zutic-2004, wang-2023, tombros-2008, quantum information shen-2022, hirohata-2025, and energy conversion and storage sendek-2018.
@@
Recent advances in ab initio theoretical methods have made it possible to reliably calculate a wide range of x-ray spectroscopies vinson-2011, apra-2020. Despite these successes, such calculations remain computationally demanding and are often limited by system size and the need to treat electron correlation pople-1999. By contrast, a trained machine learning model can make predictions orders of magnitude faster, enabling rapid exploration of material spaces.
@@
The relationship between a material’s local chemical and geometric environment, and its polarization-dependent spectral response, is inherently nonlinear and highly complex. Spectral anisotropy arises from directional variations in the electronic states, which are influenced by factors such as local bonding asymmetry, ligand field distortions, coordination number, and electronic hybridization Drager.Brummer1984. Traditional linear or descriptor-based approaches struggle to capture these coupled, high-dimensional effects. Machine learning models, by contrast, are well-suited to learn nonlinear mappings from structural and chemical features to spectral response.
@@
Machine learning has successfully been applied to several modalities of spectroscopy including x-ray absorption near-edge structure (XANESNet) rankine-2020, extended x-ray absorption fine structure timoshenko-2018, and vibrational spectroscopy gastegger-2017. The broader field of materials sciences has also demonstrated that machine learning models can successfully predict bulk material properties, such as utilizing a message-passing graph neural network (MPGNN) to capture anisotropic properties such as elastic wen-2024 and dielectric tensors lou-2024. However, these GNN approaches can be limited by the scarcity of high-quality training data and, like most neural network-based methods, suffer from a black-box nature that limits interpretability.
@@
In contrast, random forest models have demonstrated strong performance in materials spectroscopy; for example, Zheng et al. successfully applied them to identify coordination environments in XANES spectra Zheng.etal2020. Random forests can capture nonlinear and interacting effects among structural and chemical features, making them well-suited for predicting complex spectral responses while maintaining interpretability. To the authors’ knowledge, anisotropy in the VtC-XES regime remains largely unexplored and has not yet been the subject of any machine learning applications. In this paper, we develop a random forest regression (RFR) model trained on structural and chemical descriptors of 3dd transition metal complexes to predict the spectral anisotropy in VtC-XES. To quantify the degree of anisotropy, we introduce a new metric termed the spectral anisotropy matrix sum (SAMS).
@@
The rest of this article is organized as follows. In Section 2 we described the procedure for quantifying the spectral anisotropy and how we extract input features as our predictors in the model. We also address the calculation pipeline used for developing the training data set, and the procedure used to develop the random forest model. In Section 3 we present the core results of the model, namely the $R^{2}$ and mean absolute error (MAE). We also assess the relative importance of the features and compare predictive capabilities across the 10 3dd transition metals. Finally in Section 4 we discuss constraints of the model and propose future directions for how it could be improved and extended to new applications.
@@ 2.1 Spectral Anisotropy Matrix
To quantify the degree of anisotropy in a spectrum we introduce a metric based on the $L^{2}$-norm or Euclidean norm of the difference between individual polarizations bishop-2007. We start by decomposing a spectrum into the emission for three orthogonal polarization axes ($\sigma_{i},\sigma_{j},\sigma_{k}$). Calculating the square root of the integral over the difference squared, we arrive at a single number to characterize the ”difference” between two spectra. This value is normalized by the integral of the average (isotropic) spectra, $\bar{\sigma}$. We calculate this parameter for each possible pair of axes to produce the spectral anisotropy matrix (SAM) as defined in eq 1.
@@
The SAM is a zero-diagonal symmetric matrix. A visualization of the calculation process is shown in Figure 2, where the Cr VtC-XES is calculated along the $x$, $y$, and $z$ crystal axes of an example transition metal crystal system, LiCrCO3F2. The crystal structure is shown in Figure 2a, where the local structure around the Cr atoms is distorted such that there are no degenerate axes. The polarized spectra along the three crystal axes is shown in Figure 2b along with the difference curves between each pair of spectra. These are then used to calculate the Euclidean norm and build the SAM shown in Figure 2c. Taking the sum of the upper off diagonal components as shown in eq 2, we arrive at a quantitative measure, the spectral anisotropy matrix sum (SAMS). This metric captures the total amount of anisotropy in the spectrum, and is the numerical value we aim to predict with our model in this work. A SAM is highly dependent on the basis used to define the cluster around the transition metal ion. Our goal is to create a dataset of the maximum possible SAMS signal that can be achieved for a given structure, which often requires rotating to the principal axes.
@@ 2.2 Feature Anisotropy Matrices
To parameterize the anisotropy of the local atomic environment around an absorber atom, we construct 3 related but distinct metrics derived from the quadrupole moment eq 3a, dipole moment eq 3b, and inertia tensor eq 3c respectively.
@@
The quadrupole and inertia matrices are identical with the exception of the effective charge $q_{n}$, which serves to encode chemical information. We will see that the anisotropy metrics of both are strongly correlated with the spectral anisotropy, but we choose to use the eigenvectors of the weighted quadrupole moment to define a rotation-invariant basis. This is done both to maximize the observed spectral anisotropy in eq 2 and provide a set coordinate system to work in. The main reason for including both the quadrupole eq 3a and the inertia eq 3c comes from the reliance on estimating the ligand charges $q_{n}$. We rely on simple oxidation state estimations and avoid doing any direct electronic structure calculations for generating input features, and uncertainties in the charge estimates can degrade the reliability of the quadrupole metric. Incorporating a purely geometric metric thus provides a complementary, charge-independent measure of local anisotropy that can help disentangle geometric effects from uncertainties in ligand charge assignment.
The quadrupole, dipole, and inertia anisotropy matrices (denotes QAM, DAM, and IAM respectively) are shown in eq 4. They are all expressed in the eigen-basis of the quadrupole matrix (eigenvalues $\lambda$), where $\tilde{D}=V^{-1}DV$ and $\tilde{I}=V^{-1}IV$ are the dipole and inertia matrices transformed into this basis and $V$ is the matrix whose columns are the orthonormal eigenvectors of the weighted quadrupole moment.
@@
Consistent with the single anisotropy measure used for summarizing the spectral anisotropy metric into the SAMS in eq 2, a similar sum is applied to the off diagonal elements of eq 4 to arrive at the quadrupole, dipole, and inertia anisotropy matrix sums, denoted QAMS, DAMS, and IAMS respectively (eq 5). If a unit cell contains multiple unique sites for a 3dd transition metal, we average the anisotropy matrix sum over all sites to obtain a single representative number for the entire crystal.
@@ 2.3 Connection between Quadrupole and Spectra
The continuous form of the quadrupole moment tensor is defined in eq 7 (without additional distance weighting). This gives the quadrupole moment of a charge distribution with electron density $\rho_{e}(\vec{r})$ (eq 8), where $e$ is the electron charge, $N$ is the number of electrons, and $\Psi$ is the many-body wavefunction.
@@
In eq 6, the initial and final state wavefunctions are shaped by the local chemical environment. This manifests as polarization-dependent effects in the spectrum. Since the electron density in eq 8 is derived from the modulus squared of the many-body wavefunction, it retains similar spatial anisotropy, albeit without phase information. The quadrupole moment is therefore a natural option for encoding this anisotropy, capturing both geometric and electronic symmetry-breaking features of the local environment.
To further assess the relationship between the QAMS and the SAMS, we now examine how they respond to systematic geometric distortions in representative materials, specifically NiO and Cr2O3, as shown in Figure 3. Figure 3b,d show the crystal structure for NiO and Cr2O3 respectively, while Figure 3a,c show the QAMS vs. SAMS for different degrees of stretch or compression from the experimental structure. In Figure 3a the NiO structure is stretched along the $c$-axis in increments of 0.5%. Starting from the lower left-hand corner of the plot—where all bond lengths are equal—the structure is isotropic, and both QAMS and SAMS are exactly zero, which is in agreement with the perfect cubic NiO structure. As the structure becomes increasingly stretched, both QAMS and SAMS vary linearly. This suggests that for simple distortions such as bond elongation without significant angular rearrangement, QAMS tracks spectral anisotropy with high fidelity.
@@
However, in the case of a similar incremental distortion of Cr2O3 in Figure 3c we observe a distinctly nonlinear relationship between QAMS and SAMS. Here the local environment around the Cr ion is a distorted Oh structure, and has non-zero QAMS and SAMS for even the starting crystal structure. In the extremes of both compression and stretching, the relationship between the two parameters remains linear. In the region around the unchanged structure, we note that the relationship deviates from linearity, and that the minimum of the SAMS does not line up with the minimum of the QAMS.
This behavior leads to two key insights. First, the nonlinear behavior arises from changes not only in bond lengths but also in bond angles, which substantially affect the spatial charge distribution and, consequently, the spectral anisotropy. For NiO, the stretching occurs uniformly without altering angular geometry, but Cr2O3 has a more complex structural response to distortion. While the two parameters are clearly correlated, their relationship is not strictly proportional. Second, it is possible to have a local cluster that produces a zero QAMS and a non-zero SAMS (for example, a perfectly linear molecule where the ligands have the same charge $q$). The DAMS partially addresses this limitation by capturing dipolar contributions to the anisotropy, as it is derived from a metric that is odd under spatial inversion. However, this alone is insufficient, and additional features are required to more fully capture and discriminate between the underlying sources of spectral anisotropy.